# Linguistic Collapse: Neural Collapse in (Large) Language Models

### v2

@@ Abstract
Neural collapse ($\mathcal{NC}$) is a phenomenon observed in classification tasks where top-layer representations collapse into their class means, which become equinorm, equiangular and aligned with the classifiers. These behaviors – associated with generalization and robustness – would manifest under specific conditions: models are trained towards zero loss, with noise-free labels belonging to balanced classes, which do not outnumber the model’s hidden dimension. Recent studies have explored $\mathcal{NC}$ in the absence of one or more of these conditions to extend and capitalize on the associated benefits of ideal geometries. Language modeling presents a curious frontier, as training by token prediction constitutes a classification task where none of the conditions exist: the vocabulary is imbalanced and exceeds the embedding dimension; different tokens might correspond to similar contextual embeddings; and large language models (LLMs) in particular are typically only trained for a few epochs. This paper empirically investigates the impact of scaling the architectures and training of causal language models (CLMs) on their progression towards $\mathcal{NC}$. We find that $\mathcal{NC}$ properties that develop with scale (and regularization) are linked to generalization. Moreover, there is evidence of some relationship between $\mathcal{NC}$ and generalization independent of scale. Our work thereby underscores the generality of $\mathcal{NC}$ as it extends to the novel and more challenging setting of language modeling. Downstream, we seek to inspire further research on the phenomenon to deepen our understanding of LLMs – and neural networks at large – and improve existing architectures based on $\mathcal{NC}$-related properties. Our code is hosted on GitHub: https://github.com/rhubarbwu/linguistic-collapse.
Figure 1: Simultaneous development of the four neural collapse ($\mathcal{NC}$) [1] properties in 230 causal language models trained on TinyStories [2], alongside improvement in generalization (i.e. validation performance). Left to right: $\mathcal{NC}_{1}$) within-class (representation) variability collapse; $\mathcal{GNC}_{2}$) hyperspherical uniformity of class means; $\mathcal{UNC}_{3}$) uniform duality between class means and corresponding classifiers; and $\mathcal{NC}_{4}$) agreement between token (maximum a prior) classifiers and implicit nearest-class center classifiers. Colored by model size and annotated with coefficient of determination ($R^{2}$).
@@ 1.1 Neural Collapse
    Convergence to a simplex ETF: Class means tend towards equinorm and equiangular vectors when centered about the global average. The resulting geometry – known as a simplex equiangular tight frame (ETF) – maximizes pairwise angles and distances.
@@
    Nearest decision rule: Linear classifiers approximate a nearest-class center (NCC) classifier: top-layer representations predict the class with the closest mean (implied by $\mathcal{NC}$1-3).
These behaviors, often associated with improved generalization and robustness [3, 4, 5] (among other benefits, such as those discussed in §1.4), traditionally manifest under the following conditions:
@@ 4. Rq4)
$\mathcal{NC}$ is a phenomenon observed specifically in classification tasks. While not traditionally thought of as classifiers, language models – including large language models (LLMs) – learn to model aleatoric uncertainty, which can be viewed as stochastic token prediction [6]. For instance, masked language models (MLMs) such as BERT [7] predict one or several masked tokens within an input sequence based on the surrounding context. Likewise, autoregressive or causal language models (CLMs) such as GPT [8] predict the next token in a sequence given the context of previous tokens. Most of these models are essentially (pre-)trained by token classification on their vocabularies. This parallel – also drawn by [9] – raises a few natural questions:
@@ 1. $\neg$¬Rq1)
    Many classes: The unique tokens found in language modeling vocabularies are vast, usually numbering in the tens of thousands and far exceeding the hidden dimension [10].
@@ 1.4 Significance
Recently, methods building on $\mathcal{NC}$ have found use in deep learning at large. We highlight areas such as federated learning [20], graph neural networks [21], incremental/continual learning [22, 23, 24], meta-learning [24, 25], out-of-distribution detection [26, 27, 28], privacy [29, 30], learning theory [31, 32, 33, 34, 35, 36] transfer learning [3, 5, 37, 38, 39, 40], and even LLM prompting [41]. With our results, we aim to extend insights from such contributions and other related works to the autoregressive language modeling domain and ultimately assist researchers in improving and interpreting their (L)LMs.
@@ 2 Related Works
$\mathcal{NC}$ established the simplex ETF as an optimal configuration. However, a perfect simplex ETF ($\mathcal{NC}_{2}$) requires that the number of classes $C$ not exceed $d+1$ where $d$ is the embedding dimension. This requirement that $d$ be sufficiently large is not practical when the classes number beyond the thousands. [^2] Following [67], one might describe such a setting ($C>d+1$) as a model “in superposition”. For instance, GPT-2 [68] and Llama 3.1 [69] have vocabularies of around 50K and 128K tokens, respectively.
In such a scenario, one might still expect class means to be uniformly distributed on a $d$-dimensional hypersphere [18]. [19] formalize this as hyperspherical uniformity within a generalized neural collapse ($\mathcal{GNC}$) framework, which [9] then empirically demonstrate. These latter two works mention language modeling as applicable for $\mathcal{GNC}$; [9] even framed token prediction as a classification task, just as we do. We remark however that both earlier works simulate a large number of classes by drastically shrinking the embedding dimension. In contrast, we study next-token prediction, using the full class space (vocabulary) with an imbalanced token distribution and ambiguous samples.
@@
An earlier study reported that the ratio of within-class to between-class covariances of word embeddings increases with model depth [75, 76], seemingly at odds with $\mathcal{NC}_{1}$. It can however be distinguished from literature studying layer-wise $\mathcal{NC}$ in that it does not center the mean representation vectors (i.e. subtract the global mean).
[19] fine-tuned BERT [7] using their hyperspherical uniformity gap (HUG) objective on binary classification tasks from the GLUE benchmark [77]. [78] conducted a tangentially-related investigation of convolutional neural networks for few-class semi-supervised clustering in which they identify $\mathcal{NC}$. Our work is distinct from these in several ways: a) our class space far exceeds our embedding dimension ($C\gg d+1$) because we classify on the full token vocabulary; b) we analyze embeddings at a token-level granularity rather than at the sequence level; and c) our next token prediction is causal (context-dependent) as opposed to the per-sample independence of their few-category classification.
A more related work studied feature collapse in individual word representations [79], but the authors note that their analysis is limited to shallow NLP modeling on more rigid and tabular text data.
@@ 3.1 Dataset and Language Models
TinyStories [2] is a synthetic444 TinyStories was developed and evaluated as a faithful emulation of large language modeling, so we chose it for experimentation to train hundreds of CLMs and analyze embeddings at low cost. See Appendix A. dataset generated by GPT-3.5 and GPT-4 using around 1500 English words a child might use. Next-token prediction is performed by sampling from the token vocabulary $\mathbb{V}=\llbracket 1,29233\rrbracket$, which for our purposes can therefore be framed as classification among $C=$29,233$$ classes. [^5] Although the GPT-Neo [80] tokenizer has over 50K tokens, only a subset vocabulary appears in TinyStories. Following the GPT-style preprocessing regime [8], raw sequences are packed into $S$ chunks of size $T$, providing $N=S(T-1)$ token samples for training. [^6] We cannot use the first ground-truth nor last predicted token in any chunk. Details are listed in Appendix A.
@@ 3.3 Homogeneity and Variability
3.4 Signal-to-Noise Ratio – $\mathcal{NC}1$
The ability to disambiguate between classes depends on the ratio of within-class to between-class variabilities. Building upon foundational works [85, 86], $\mathcal{NC}$ originally measured variability through an inverse signal-to-noise ratio (SNR), whose minimization constitutes within-class variability collapse ($\mathcal{NC}_{1}$). We instead employ a class-distance normalized variance (CDNV) similar to [3]:
@@ 3.4 Signal-to-Noise Ratio – $\mathcal{NC}1$
3.5 Geometric Structures – $\mathcal{(G)NC}2$
@@
$\mathcal{NC}_{2}$ also entails that class means are equiangular about their center $\boldsymbol{\bar{\mu}}$: pairwise distances and angles between their class means should be maximized and similar. Following [1], we measure interference (sometimes known as similarity or coherence [89, 90]). Its minimization,
@@
3.6 Alignment Between Classifiers and Class Mean Embeddings – $\mathcal{(U)NC}3$
@@
3.7 Agreement of the Classifiers – $\mathcal{NC}4$
Finally, $\mathcal{NC}_{4}$ is described as the simplification of the linear classifier’s MAP predictions (Equation 1) for test [^15] We used the validation set since it was not used for training or hyperparameter optimization. points to those of the implicit nearest-class center (NCC) classifier:
@@ 4 Experimental Results
4.1 Within-Class Variability Collapse – $\mathcal{NC}1$
Scaling our models dramatically reduces normalized variance, which is further aided by more training epochs and stronger weight decay (Appendix Figs. 6, 7). These noise reductions tightly associate with generalization (Fig. 1, left, “$\mathcal{NC}_{1}$”). The relationship is most apparent at larger model scale.
4.2 Geometric Structures – $\mathcal{(G)NC}2$
Figure 2: Validation loss shows negligible correlation with equinormness ($\mathcal{NC}_{2}$, left), some relationship with equiangularity ($\mathcal{NC}_{2}$, centre), and a stronger one with hyperspherical uniformity ($\mathcal{GNC}_{2}$, right). So, $\mathcal{GNC}_{2}$ develops with scale and correlates well with generalization, better than $\mathcal{NC}_{2}$.
@@ 4.2 Geometric Structures – $\mathcal{(G)NC}2$
Class mean embedding norms grow with model width and training (Appendix Fig. 8), and subtly with depth (Appendix Fig. 9). Despite this growth, the variation of norms consistently decreases (improving equinormness) with scale (Appendix Figs. 10, 11). Both trends correlate with improved generalization (Appendix Fig. 20).
@@
Scaling model dimensions reduces average interference (Appendix Figs. 12, 13) down to an empirical plateau of approximately $10^{-3}$, which is more apparent in less-trained models. However, the variation of interference rises and persists when scaling (Appendix Figs. 14, 15), suggesting that interference becomes more concentrated between some pairs of classes. These results could be due to various factors, including but not limited to unfriendly conditions of language modeling (§1.2) or the impossibility of a simplex ETF (§3.5).
We only see a rough correlation between average interference and generalization and almost none with its variation, equiangularity (Appendix Fig. 21). In other words, the limited trends we observed toward a simplex ETF do not appear to be associated with generalization when $C>d+1$.
@@
4.3 Classifier (Mis)alignment and Duality – $\mathcal{(U)NC}3$
Model width ($d$) is correlated faintly with average similarity between class means and their respective classifiers (Appendix Fig. 23), but strongly with variation in similarity (Appendix Fig. 25). The relationships to generalization follow the same pattern (Fig. 3), suggesting that uniform duality ($\mathcal{UNC}_{3}$) might serve as a better $\mathcal{NC}$ property than self-duality ($\mathcal{NC}_{3}$) overall; we discuss this in §5.1.
@@ 4.3 Classifier (Mis)alignment and Duality – $\mathcal{(U)NC}3$
4.4 Classifier Agreement – $\mathcal{NC}4$
The linear and NCC classifiers agree on far more examples than random chance. Scaling encourages agreement (Appendix Figs. 29, 30). Increasing width for certain depths happens to plateau or even regress the agreement rate, but this limitation is overcome with further training. And finally, agreement is a strong indicator of generalization (Fig. 1, right, $\mathcal{NC}_{4}$).
@@ 5.2 Duality of Duality
The dichotomy of self-duality ($\mathcal{NC}_{3}$) and uniform duality ($\mathcal{UNC}_{3}$) is rather conspicuous. Our main experiments find that $\mathcal{NC}_{3}$ does not consistently develop with scale while $\mathcal{UNC}_{3}$ does (Fig. 3). However, within fixed scale, the opposite is true, implying that $\mathcal{UNC}_{3}$ may be confounded by model capacity while $\mathcal{NC}_{3}$ is a subtle and fine-grained indicator of generalization.
@@ 5.3 The Effect of Weight Regularization
Our models trained with either weight decay factor exhibited very similar patterns in the emergence of $\mathcal{NC}$ (or lack thereof), but the more aggressive factor $\beta=0.1$ resulted in stronger development of $\mathcal{NC}$ properties than with $\beta=0.0005$ (Appendices E, F, G, H, I, J, K, M, N, P). These findings empirically affirm $\beta=0.1$ weight decay as sensible for LLM pre-training [81], and concur with [56] on the pivotal role that appropriate regularization plays in the emergence of $\mathcal{NC}$.

### v3

@@ Abstract
Neural collapse ($\mathcal{NC}$) is a phenomenon observed in classification tasks where top-layer representations collapse into their class means, which become equinorm, equiangular and aligned with the classifiers. These behaviours — associated with generalization and robustness — would manifest under specific conditions: models are trained towards zero loss, with noise-free labels belonging to balanced classes, which do not outnumber the model’s hidden dimension. Recent studies have explored $\mathcal{NC}$ in the absence of one or more of these conditions to extend and capitalize on the associated benefits of ideal geometries. Language modelling presents a curious frontier, as training by token prediction constitutes a classification task where none of the conditions exist: the vocabulary is imbalanced and exceeds the embedding dimension; different tokens might correspond to similar contextual embeddings; and large language models (LLMs) in particular are typically only trained for a few epochs. This paper empirically investigates the impact of scaling the architectures and training of causal language models (CLMs) on their progression towards $\mathcal{NC}$. We find that $\mathcal{NC}$ properties that develop with scale (and regularization) are linked to generalization. Moreover, there is evidence of some relationship between $\mathcal{NC}$ and generalization independent of scale. Our work thereby underscores the generality of $\mathcal{NC}$ as it extends to the novel and more challenging setting of language modelling. Downstream, we seek to inspire further research on the phenomenon to deepen our understanding of LLMs — and neural networks at large — and improve existing architectures based on $\mathcal{NC}$-related properties. Our code is hosted on GitHub: https://github.com/rhubarbwu/linguistic-collapse.
Figure 1: Simultaneous development of the four neural collapse ($\mathcal{NC}$) [1] properties in 230 causal language models trained on TinyStories [2], alongside improvement in generalization (i.e. validation performance). Left to right: $\mathcal{NC}_{1}$) within-class (representation) variability collapse; $\mathcal{GNC}_{2}$) hyperspherical uniformity of class means; $\mathcal{UNC}_{3}$) uniform duality between class means and corresponding classifiers; and $\mathcal{NC}_{4}$) agreement between token (maximum a prior) classifiers and implicit nearest-class centre classifiers. Coloured by model size and annotated with coefficient of determination ($R^{2}$).
@@ 1.1 Neural Collapse
    Convergence to a simplex ETF: Class means tend towards equinorm and equiangular vectors when centred about the global average. The resulting geometry — known as a simplex equiangular tight frame (ETF) — maximizes pairwise angles and distances.
@@
    Nearest decision rule: Linear classifiers approximate a nearest-class centre (NCC) classifier: top-layer representations predict the class with the closest mean (implied by $\mathcal{NC}$1-3).
These behaviours, often associated with improved generalization and robustness [3, 4, 5] (among other benefits, such as those discussed in §1.4), traditionally manifest under the following conditions:
@@ 4. Rq4)
$\mathcal{NC}$ is a phenomenon observed specifically in classification tasks. While not traditionally thought of as classifiers, language models — including large language models (LLMs) — learn to model aleatoric uncertainty, which can be viewed as stochastic token prediction [6]. For instance, masked language models (MLMs) such as BERT [7] predict one or several masked tokens within an input sequence based on the surrounding context. Likewise, autoregressive or causal language models (CLMs) such as GPT [8] predict the next token in a sequence given the context of previous tokens. Most of these models are essentially (pre-)trained by token classification on their vocabularies. This parallel — also drawn by [9] — raises a few natural questions:
@@ 1. $\neg$¬Rq1)
    Many classes: The unique tokens found in language modelling vocabularies are vast, usually numbering in the tens of thousands and far exceeding the hidden dimension [10].
@@ 1.4 Significance
Recently, methods building on $\mathcal{NC}$ have found use in deep learning at large. We highlight areas such as federated learning [20], graph neural networks [21], incremental/continual learning [22, 23, 24], meta-learning [24, 25], out-of-distribution detection [26, 27, 28], privacy [29, 30], learning theory [31, 32, 33, 34, 35, 36] transfer learning [3, 5, 37, 38, 39, 40], and even LLM prompting [41]. With our results, we aim to extend insights from such contributions and other related works to the autoregressive language modelling domain and ultimately assist researchers in improving and interpreting their (L)LMs.
@@ 2 Related Works
$\mathcal{NC}$ established the simplex ETF as an optimal configuration. However, a perfect simplex ETF ($\mathcal{NC}_{2}$) requires that the number of classes $C$ not exceed $d+1$ where $d$ is the embedding dimension. This requirement that $d$ be sufficiently large is impractical when the classes number beyond the thousands. [^2] Following [67], one might describe such a setting ($C>d+1$) as a model “in superposition”. For instance, GPT-2 [68] and Llama 3.1 [69] have vocabularies of around 50K and 128K tokens, respectively.
In such a scenario, one might still expect class means to be uniformly distributed on a $d$-dimensional hypersphere [18]. [19] formalize this as hyperspherical uniformity within a generalized neural collapse ($\mathcal{GNC}$) framework, which [9] then empirically demonstrate. These latter two works mention language modelling as applicable for $\mathcal{GNC}$; [9] even framed token prediction as a classification task, just as we do. We remark however that both earlier works simulate a large number of classes by drastically shrinking the embedding dimension. In contrast, we study next-token prediction, using the full class space (vocabulary) with an imbalanced token distribution and ambiguous samples.
@@
An earlier study reported that the ratio of within-class to between-class covariances of word embeddings increases with model depth [75, 76], seemingly at odds with $\mathcal{NC}_{1}$. It can, however, be distinguished from literature studying layer-wise $\mathcal{NC}$ in that it does not centre the mean representation vectors (i.e. subtract the global mean).
[19] fine-tuned BERT [7] using their hyperspherical uniformity gap (HUG) objective on binary classification tasks from the GLUE benchmark [77]. [78] conducted a tangentially-related investigation of convolutional neural networks for few-class semi-supervised clustering in which they identify $\mathcal{NC}$. Our work is distinct from these in several ways: a) our class space far exceeds our embedding dimension ($C\gg d+1$) because we classify on the full token vocabulary; b) we analyze embeddings at a token-level granularity rather than at the sequence level; and c) our next-token prediction is causal (context-dependent) as opposed to the per-sample independence of their few-category classification.
A more related work studied feature collapse in individual word representations [79], but the authors note that their analysis is limited to shallow NLP modelling on more rigid and tabular text data.
@@ 3.1 Dataset and Language Models
TinyStories [2] is a synthetic444 TinyStories was developed and evaluated as a faithful emulation of large language modelling, so we chose it for experimentation to train hundreds of CLMs and analyze embeddings at a low cost. See Appendix A. dataset generated by GPT-3.5 and GPT-4 using around 1500 English words a child might use. Next-token prediction is performed by sampling from the token vocabulary $\mathbb{V}=\llbracket 1,29233\rrbracket$, which for our purposes can therefore be framed as classification among $C=$29,233$$ classes. [^5] Although the GPT-Neo [80] tokenizer has over 50K tokens, only a subset vocabulary appears in TinyStories. Following the GPT-style preprocessing regime [8], raw sequences are packed into $S$ chunks of size $T$, providing $N=S(T-1)$ token samples for training. [^6] We cannot use the first ground-truth nor the last predicted token in any chunk. Details are listed in Appendix A.
@@ 3.3 Homogeneity and Variability
3.4 Signal-to-Noise Ratio — $\mathcal{NC}1$
The ability to disambiguate between classes depends on the ratio of within-class to between-class variabilities. Building upon foundational works [85, 86], $\mathcal{NC}$ originally measured variability through an inverse signal-to-no ratio (SNR), whose minimization constitutes within-class variability collapse ($\mathcal{NC}_{1}$). We instead employ a class-distance normalized variance (CDNV) similar to [3]:
@@ 3.4 Signal-to-Noise Ratio — $\mathcal{NC}1$
3.5 Geometric Structures — $\mathcal{(G)NC}2$
@@
$\mathcal{NC}_{2}$ also entails that class means are equiangular about their centre $\boldsymbol{\bar{\mu}}$: pairwise distances and angles between their class means should be maximized and similar. Following [1], we measure interference (sometimes known as similarity or coherence [89, 90]). Its minimization,
@@
3.6 Alignment Between Classifiers and Class Mean Embeddings — $\mathcal{(U)NC}3$
@@
3.7 Agreement of the Classifiers — $\mathcal{NC}4$
Finally, $\mathcal{NC}_{4}$ is described as the simplification of the linear classifier’s MAP predictions (Equation 1) for test [^15] We used the validation set since it was not used for training or hyperparameter optimization. points to those of the implicit nearest-class centre (NCC) classifier:
@@ 4 Experimental Results
4.1 Within-Class Variability Collapse — $\mathcal{NC}1$
Scaling our models dramatically reduces normalized variance, which is further aided by more training epochs and stronger weight decay (Appendix Figs. 6, 7). These noise reductions tightly associate with generalization (Fig. 1, left, “$\mathcal{NC}_{1}$”). The relationship is most apparent at scale.
4.2 Geometric Structures — $\mathcal{(G)NC}2$
Figure 2: Validation loss is correlated with all three measurements: (left) equinormness ($\mathcal{NC}2$) expressed as variation in logarithmic norms; (centre) equiangularity ($\mathcal{NC}2$) as variation in interference; (right) hyperspherical uniformity ($\mathcal{GNC}2$) as variation in logarithmic pairwise distances.
@@ 4.2 Geometric Structures — $\mathcal{(G)NC}2$
Logarithmic class mean norms grow with model width and training (Appendix Fig. 8), and subtly with depth (Appendix Fig. 9). Meanwhile, the variation of these (logarithmic) norms consistently decreases (improving equinormness) with scale (Appendix Figs. 10, 11). Both trends correlate with improved generalization (Appendix Fig. 20).
@@
Scaling model dimensions reduces average interference (Appendix Figs. 12, 13) down to an empirical plateau of approximately $10^{-3}$, which is more apparent in less-trained models. However, the variation of interference rises and persists when scaling (Appendix Figs. 14, 15), suggesting that interference becomes more concentrated between some pairs of classes. These results could be due to various factors, including but not limited to unfriendly conditions of language modelling (§1.2) or the impossibility of a perfect simplex ETF (§3.5).
Appendix Figure 21 shows only a rough performance correlation with average interference (i.e. coherence) and a more definite — albeit still noisy — one with the variation of interference (i.e. equiangularity). In other words, the limited trends we observed toward a simplex ETF suggest that the association of $\mathcal{NC}2$ with generalization begins to break down when $C>d+1$.
@@
4.3 Classifier (Mis)alignment and Duality — $\mathcal{(U)NC}3$
Model width ($d$) is correlated faintly with the average similarity between class means and their respective classifiers (Appendix Fig. 23), but strongly with variation in similarity (Appendix Fig. 25). The relationships to generalization follow the same pattern (Fig. 3), suggesting that uniform duality ($\mathcal{UNC}_{3}$) might serve as a better $\mathcal{NC}$ property than self-duality ($\mathcal{NC}_{3}$) overall; we discuss this in §5.1.
@@ 4.3 Classifier (Mis)alignment and Duality — $\mathcal{(U)NC}3$
4.4 Classifier Agreement — $\mathcal{NC}4$
The linear and NCC classifiers agree on far more examples than random chance, and model scaling encourages this agreement (Appendix Figs. 29, 30). Increasing width for certain depths happens to plateau or even regress the agreement rate, but this limitation is overcome with further training. And finally, agreement is a strong indicator of generalization (Fig. 1, right, $\mathcal{NC}_{4}$).
@@ 5.2 Duality of Duality
The dichotomy of self-duality ($\mathcal{NC}_{3}$) and uniform duality ($\mathcal{UNC}_{3}$) is rather conspicuous. Our main experiments find that $\mathcal{NC}_{3}$ does not consistently develop with scale while $\mathcal{UNC}_{3}$ does (Fig. 3). However, within a fixed scale, the opposite is true, implying that $\mathcal{UNC}_{3}$ may be confounded by model capacity while $\mathcal{NC}_{3}$ is a subtle and fine-grained indicator of generalization.
@@ 5.3 The Effect of Weight Regularization
Our models trained with either weight decay factor exhibited very similar patterns in the emergence of $\mathcal{NC}$ (or lack thereof), but the more aggressive factor $\beta=0.1$ resulted in stronger development of $\mathcal{NC}$ properties than with $\beta=0.0005$ (Appendices E, F, G, H, I, J, K, M, N, P). These findings empirically affirm $\beta=0.1$ weight decay as sensible for CLM pre-training [81], and concur with [51] on the pivotal role that appropriate regularization plays in the emergence of $\mathcal{NC}$.